find the value of sin[cot inverse{cos(tan inverse 1)}]

Grade:12

1 Answers

Arun
5 years ago
 
Dear student
 
start from the most right side 
 
sin ( cot-1 (cos (tan-1 1)))
 
= sin ( cot-1 (cos pi/4))
 
= sin ( cot-1 (1/ sqrt 2))
 
 = sin (sin-1 sqrt(2/3))
 
 = sqrt (2/3)